A jogger ran a 30km race at a speed of x km/hr. He was 4 hours faster than a walker who was walking 10km/hr slower. What speed was the jogger jogging at?
I have been stuck on this question for a while now but am getting nowhere.
Could you please help?
Thank you.

Answer by mananth(16949)

Time taken by walker - time taken by jogger = 4 hours
let walker speed = x kph
jogger speed = x+4 kph
distance = 30
t=d/r
30/(x-10)-30/x=4
multiply equation by x(x-10)
30x-30(x-10)= 4x(x-10)
30x+300-30x=4x^2-40x
4x^2-40x-300=0
/4
x^2-10x-75=0
x^2-15x+10x-75=0
x(x-15)+10(x-15)=0
(x-15)(x+10)=0
x=15 OR -10
ignore negative
x=15